OpenStudy (anonymous):

please help with systems of equations! y=3x+1 4y=12x+4 (I need to find y and x)

OpenStudy (nirmalnema):

4y=4(3x+1) y=3x+1 and the other line is also y=3x+1 so both the line are same so it has infinitely many solution

OpenStudy (nirmalnema):

@tif49 you get that

OpenStudy (anonymous):

thank youu sooo muchh :)!!!

OpenStudy (anonymous):

i do, thanks !!

OpenStudy (nirmalnema):

your welcome...!!